OpenStudy (anonymous):

sinxcos(π/7)-sin(π/7)cosx=-√2/2, then x can equal: _____. Check all that apply. A. 7π/4+π/7+2nπ B. 5π/4+π/7+2nπ C. P/4+π/7+2nπ D. 3π/4+π/7+2nπ

OpenStudy (stacey):

sinx cos(π/7) - sin(π/7) cosx = sin (x - π/7) To get -√2/2, the angle must be 5π/4 + 2nπ or 7π/4 + 2nπ, so x - π/7 = 5π/4 + 2nπ or x - π/7 = 7π/4 + 2nπ

OpenStudy (stacey):

To solve for x, we must add π/7 to both sides in both of the equations above.